North Star had the following data (thousands of dollars):

Cash and equivalents

100.00

Fixed assets

283.50

Sales

1,000.00

Net income

50.00

Current liabilities

105.50

Notes payable to bank

20.00

Current ratio

3.00

DSO

40.55 days

ROE

12.00%

North Star has no preferred stock—only common equity, current liabilities, and long-term debt.

a.Find North Star’s (2) accounts receivable.

b.Find North Star’s (3) current assets.

c.Find North Star’s (4) total assets.

d.Find North Star’s (6) common equity.

e.Find North Star’s (7) inventory.

f.Find North Star’s (9) long-term debt.
Answer #1

a. Accounts Receivable: $ 111.10 thousand.

DSO = (365 / Sales) * Accounts Receivable = 40.55

or 365 / 1,000 * Accounts Receivable = 40.55

Accounts Receivable = 40.55 / 0.365 = 111.10

b. Current Asset: $ 316.50 thousand

Current Ratio = Current Assets / Current Liabilities = Current Assets / 105.50 = 3

or Current Assets = 105.50 x 3 = 316.50

c. Total Assets = $ 600 thousand.

Total Assets = Fixed Assets + Current Assets = $ 283.50 + $ 316.50 = 600

d. Common Equity : $ 416.67 thousand.

Return on Equity = Net Income / Common Equity = 50 / Common Equity = 0.12

or Common Equity = 50 / 0.12 = 416.67

e. Inventory : $ 105.4 thousand.

f. Long term debt : $ 57.83 thousand.